# Trajectory Balance with Asynchrony: Decoupling Exploration and Learning for Fast, Scalable LLM Post-Training

### v1

@@ 4 TBA: Fast, Scalable LLM Post-Training
TBA is an asynchronous distributed RL framework for post-training language models (see Figure 2 for a visualization). TBA uses the off-policy trajectory balance objective (Equation 5) to efficiently leverage scaled data generation, and it decouples data generation from model updates to ensure high resource utilization. TBA has two key components: a single Trainer node and one or more Searcher nodes that collect off-policy data into a shared replay buffer $\mathcal{D}_{\mathrm{global}}$.
TBA can be implemented by, e.g., modifying the RLOO trainer class of the Hugging Face TRL library (von Werra et al., 2020) to use the TB objective, creating a dictionary to hold trajectories, and assigning training and search to distinct nodes. Here, a node is a computational resource sufficient to perform all the operations needed for training or search – it can be a collection of multiple GPUs or even a subset of 1 GPU. In our experiments, a node is always 1 GPU; e.g., given 16 GPUs, we would have 15 Searcher nodes and 1 Trainer node operating.
Separation of the Searcher and Trainer is highly desirable even in a 2 node (e.g. 2 GPU) cluster because LLM policy rollouts are costly sequential decoding procedures, and training requires only parallel likelihood evaluation of an entire sequence through a single forward pass. Thus, given an objective that can tolerate off-policy asynchronous online training, massive wall clock time speedups can be realized by continuously running training on 1 node without pausing for rollout generation.
@@ 4.1 Scaling data collection with Searcher nodes
In TBA, the Searcher nodes each carry a local delayed copy $\pi_{\theta^{\prime}}$ of the Trainer policy $\pi_{\theta}$. To produce policy rollouts, queries $\mathbf{x}$ are sampled from a dataset, and the local policy generates a batch of $K$ responses $\mathbf{y}\sim\pi_{\theta^{\prime}}(\mathbf{y}|\mathbf{x})$ that are evaluated with the reward model $r_{\phi}(\mathbf{y};\mathbf{x})$. Like Noukhovitch et al. (2025), we use vLLM (Kwon et al., 2023) for faster generation. The $(\mathbf{x},\mathbf{y},r_{\phi}(\mathbf{y};\mathbf{x}))$ tuples are stashed in the Searcher’s local replay buffer $\mathcal{D}_{\text{local}}$ (note that $\mathbf{x}$ can instead be stored as a dictionary key to save space). We also add to the stored tuple the step of the trainer when syncing last occurred, giving us a notion of how off-policy the generated data is – this can later be used by the Trainer to prioritize sampling from more recent generations for relatively “on-policy” updates.
Periodically, search and training pause to pull each Searcher node’s local replay buffer $\mathcal{D}_{\text{local}}$ into the global replay buffer $\mathcal{D}_{\mathrm{global}}$, and to update the searcher’s local policy with the trainer’s version. The global buffer maintains a list of all generated responses and rewards for each query.
A key motivation for scaling up data generation through a large number of Searcher nodes is improving exploration. For example, we generate $S>K$ samples for a given query, even when only updating the model using $K$ samples per query – this can mitigate the lack of diversity caused by the fact that $K$ independent model rollouts are not guaranteed to produce $K$ unique sequences (when duplicates are encountered, we keep the most recently generated version). Relatedly, future work could apply simple off-policy inference techniques in the Searcher nodes such as randomly sampling the softmax temperature, or using alternative decoding techniques like beam search. We expect such approaches to particularly aid solution discovery in sparse reward settings.
@@ 5.1 Tasks
  • Mathematical reasoning (MR): We study the GSM8K task which consists of grade-school level math problems and a binary reward based on exact match for the correct final answer (Cobbe et al., 2021). We adopt the setup from Kazemnejad et al. (2024); Noukhovitch et al. (2025), using an SFTed RhoMath-1B (Lin et al., 2024) model as a base for RL post-training.
  • Preference fine-tuning (PFT): We consider the task of fine-tuning a language model with a reward function learned from human preference data. Specifically, we study the TL;DR summarization task where the goal is to write short summaries for reddit posts (Stiennon et al., 2020). Following Noukhovitch et al. (2025), we consider Pythia (Biderman et al., 2023) as the base model for the policy and the reward models, using the SFTed versions used by Noukhovitch et al. (2025).
  • Red-teaming (RT): We investigate automated red-teaming, another critical step for LLM post-training. The goal is to discover prompts that elicit harmful responses from a target model, as measured by a toxicity classifier. We follow the setup from Lee et al. (2025), applying the same models: our smaller-scale experiments use GPT-2 (Radford et al., 2019) as an attacker model, GPT-2 (instruction-tuned) as a victim model, and a RoBERTa-based toxicity classifier (Vidgen et al., 2021); our larger-scale experiments use Llama-3.2-1B (Meta AI Llama Team, 2024) as an attacker model, Llama-3.1-8B-Instruct as a victim model, and LlamaGuard-3-8B (Meta AI Llama Team, 2024) for measuring toxicity, averaging over multiple responses from the victim model.
For all tasks, we study solely the RL post-training components of baselines’ workflows. In particular, we start with SFTed checkpoints, then perform RL. Notably, Lee et al. (2025) also include a maximum likelihood estimation training phase after RL that uses the buffer produced during RL, which boosts performance but is not investigated here. Using these baselines’ codes, we follow their setup and hyperparameters, with deviations noted in Appendix A. For MR and PFT, we implement TBA by augmenting the RLOO trainer of Noukhovitch et al. (2025) with our distributed asynchronous RL framework and the TB objective (Equation 5). For RT, we implement TBA by augmenting the trajectory balance trainer of Lee et al. (2025) with our distributed asynchronous RL framework.
@@ 5.2 Metrics and baselines
MR Metrics. We follow the evaluation setup of Noukhovitch et al. (2025), computing the pass@1 on the GSM8K (Cobbe et al., 2021) test dataset with greedy decoding.
MR Baselines. Prior work (Kazemnejad et al., 2024; Noukhovitch et al., 2025) applies RL post-training with the GSM8K training set to the SFTed RhoMath-1B (Lin et al., 2024) model, which initially obtains 40.3% accuracy on the test set. We compare TBA post-training to the methods used in these prior works: VinePPO (Kazemnejad et al., 2024), Online-DPO (Guo et al., 2024), PPO (Schulman et al., 2017), and RLOO (Ahmadian et al., 2024).
PFT Metrics. We follow the evaluation setup of Noukhovitch et al. (2025), using the win-rate under a 6.7B “gold” reward model (Huang et al., 2024) as the primary metric. We additionally report approximate KL distance—approximated by perplexity to match the evaluation of Noukhovitch et al. (2025)—between the learned policy and the reference policy.
PFT Baselines. Following Noukhovitch et al. (2025), we compare TBA with: Online-DPO (Guo et al., 2024), PPO (Schulman et al., 2017), and RLOO (Ahmadian et al., 2024).
RT Metrics. We follow Lee et al. (2025), measuring the attack success rate on 1024 sampled prompts for a victim model. We also measure the diversity of these test-time generated prompts by computing the average pairwise cosine distance.
RT Baselines. We compare against: SFT, PPO (Schulman et al., 2017) (with the novelty reward from Lee et al. (2025)), REINFORCE (Williams, 1992; Sutton et al., 1999), RLOO (Ahmadi & Mahmudi, 2023), Online DPO (Rafailov et al., 2023), GFlowNet (one-actor TB) (Lee et al., 2025).
Figure 3: TBA scales search and improves RL efficiency on the TL;DR summarization task. All plotted points use 4xA100 GPUs, but TBA allocates 3 GPUs to search, and Online DPO allocates 1 GPU to search. TBA produces large-scale off-policy data that its trajectory balance objective can leverage, creating massive efficiency benefits. Online DPO baselines taken from Noukhovitch et al. (2025). Dashed and solid lines use 256 and 425 updates, respectively. Appendix A has details.
Figure 4: TBA defines a new KL vs. win-rate Pareto frontier for the TL;DR summarization task. The baseline “Online DPO” frontier is created by increasing the degree of off-policyness, starting from on-policy Online DPO, results from (Noukhovitch et al., 2025). The TBA frontier is created by altering the training steps, searcher count, and KL annealing schedule as described in Appendix A.
Table 1: TBA Pareto dominates baseline on-policy algorithms in the PFT task (Pythia 410M). Baseline results taken from Noukhovitch et al. (2025). Please see Appendix A for details.
Method Perplexity/KL $\downarrow$ Win Rate $\uparrow$
Online DPO 1.13 0.85
PPO 1.14 0.86
RLOO 1.13 0.82
TBA (Ours) 1.13 0.86
@@ 5.3 TBA redefines efficiency-performance Pareto frontiers
When tested on established RL problems, an alternative possibility is that TBA will underperform due to its departures from conventional approaches: it is asynchronous, off-policy, reliant on the less-common trajectory balance objective, and makes updates using many responses per query. [^1] We compute the TBA loss with more samples per query than analogous approaches like RLOO (e.g., 20 samples versus 4) to reduce the variance of the gradient of the TB objective estimate (Equation 5) that we optimize. We use fewer queries per batch to keep the batch size small despite this response scaling. Indeed, Noukhovitch et al. (2025) contemporaneously suggests potential limitations to asynchronous, off-policy RL for LLMs, finding that increasing off-policyness can harm performance metrics like win-rate or exacerbate policy deviations from the reference model. Further, Hou et al. (2024) found limited benefits to scaling response generation from 4 to 8 (or 16) responses when using PPO.
@@
We study this question by computing Pareto frontiers for MR (Figure 1), for PFT (Figures 3 and 4 and Table 1), and for RT (Figure 5 and Table 5.4). See Appendix A for experimental details. Notably, TBA produces results on or beyond the Pareto frontiers of all three tasks at multiple model scales, consistent with our hypothesis that TBA can efficiently train LLMs on off-policy data.
@@
Speed is vastly improved with TBA training, which proceeds entirely asynchronously without training-bound or generation-bound processes – the only non-training time occurs briefly every $k$ steps. In the compute-matched MR experiments (Figure 1), TBA speeds up the training of the only method with comparable performance (VinePPO) by nearly 50×50$\times$50 ×, while improving accuracy by $1.8$% and speed by 1.5×1.5$\times$1.5 × relative to the speed-optimized asynchronous DPO baseline (Noukhovitch et al., 2025). In the compute-matched PFT experiments (Figure 3), TBA produces 5$\approx$ 5$\times$ 5 × speedups over speed-optimized asynchronous DPO baselines (Noukhovitch et al., 2025). In the non-compute-matched automated red-teaming experiments (Section 5.4), TBA gives 7$\approx$ 7$\times$ 7 × speedups for GPT-2 and Llama 3.2 1B compared to the non-distributed, synchronous GFlowNet baseline (Lee et al., 2025). these results suggest that TBA is an effective, parallel, and scalable search framework for distributed learning, offering substantial speed-ups while remaining competitive with leading approaches.
@@ 5.4 Does off-policyness hurt performance?
The results in the prior section are perhaps surprising given evidence of off-policyness’s harmfulness to RL post-training (Noukhovitch et al., 2025). Thus, we investigate the effect of off-policyness with TBA by studying its related hyperparameters, which we first review here. The fraction $m$ controls how often sampling attempts to approximate an “on-policy” distribution. When $m=1$, training occurs exclusively with samples added in the most recent sync step, while $m=0$ corresponds to selecting data without regard for how recently it was generated (e.g., using reward weighting). Additionally, recall that parameters and buffer data are shared between searchers and trainers every $k$ training steps: since TBA runs exploration and training in parallel, TBA trains off-policy even when $m=1$ and $k=1$. Specifically, with probability $m$, TBA selects a sample from the central replay buffer that is at most $12k-$ 1 updates off-policy. With probability $1-m$, TBA samples data produced by the model at any point, which can be as off-policy as the number of training steps minus $1$.
@@
Importantly, the choice of reinforcement learning (RL) algorithm is crucial in determining how effectively we can leverage off-policy updates. TB is a fully off-policy compatible algorithm, and as shown in Figure 4, it significantly outperforms asynchronous Online DPO, even in the latter’s most on-policy setting. Noukhovitch et al. (2025) identified Online DPO as the best-performing off-policy algorithm in their experiments, making it particularly insightful that TB improves upon this.
Figure 5: TBA reaches the RT diversity-toxicity Pareto frontier and improves as search is scaled. (Left) On the GPT-2 automated red-teaming task of Lee et al. (2025), TBA produces results on the diversity vs. toxicity Pareto frontier in less training time. Baselines taken from Lee et al. (2025). (Right) Each searcher uses one V100 GPU for generating attacks. We report means and standard errors from multiple runs of the automated red-teaming task with GPT-2 at each searcher/GPU count.
Table 2: TBA speeds up the wall-clock time required to reach the Pareto frontier for the red-teaming task. The GFlowNet performances are taken from Lee et al. (2025), while the training speeds are computed by us with their code. With the GPT-2 models, TBA performance improves with searcher count. With the Llama models, we trade attack toxicity for attack diversity by scaling the TBA buffer’s maximum size from 130,000 to 150,000 samples*, retaining more off-policy data.
{widetable}
lllrlll Attacker / Training Method Hardware Time (h) Speedup Cosine % Toxic
Victim Model Distance Prompts
GPT-2 / GPT-2 + SFT
GFlowNet - Sync 1$\times$×V100 11.9 1x 0.65 96.6
TBA (Ours) 4$\times$×V100 1.7 7x 0.42 94.5
TBA (Ours) 16$\times$×V100 2.5 4.8x 0.45 96
TBA (Ours) 64$\times$×V100 2.9 4.1x 0.49 97.6
Llama 3.2 1B / Llama 3.1 8B - Instruct GFlowNet - Sync 1$\times$×A100 37.4 1x 0.32 100.0
TBA (Ours) 8$\times$×A100 5.7 6.6x 0.35 98.1
TBA (Ours)* 8$\times$×A100 5.7 6.6x 0.37 94.8

### v2

@@ 4 TBA: Fast, Scalable LLM Post-Training
TBA integrates the TB gradient of Eq. 5 into an asynchronous distributed RL framework for post-training language models. Asynchronous RL decouples data generation from model updates, enhancing resource utilization and reducing training duration. At the same time, the off-policy TB objective promotes efficient learning from the off-policy training data induced by asynchrony.
We test two TBA variants. The first implements TBA from scratch by modifying an RLOO trainer class (e.g. of the Hugging Face TRL library [vonwerra2022trl]) to use Eq. 5, adding a schedule for $\beta$, integrating a replay buffer and sampling approach, and parallelizing training and search to allow async RL – we discuss the details of this variant below. The second variant TBA uses the asynchronous PRIME-RL codebase [primeintellect2025primeRL], to which we simply add the update rule corresponding to Eq. 7 and reference-policy resetting. TBA has fewer tunable knobs than TBA and supports multi-GPU training processes, allowing us to test TB for async RL of LLMs in a simpler setting and with larger models/contexts.
While TBA prioritizes ease of use and multi-GPU training, our from-scratch TBA implementation prioritizes speed and search. For example, ensuring the data has a constant level of off-policyness (as PRIME-RL does) can reintroduce the bottleneck that async RL aims to remove, so TBA runs the training and searching processes continuously and independently, only syncing every $k$ steps (the “sync period”). Moreover, every completion, regardless of how stale it is, is accessible in TBA’s replay buffer (e.g. via reward-based sampling). As discussed below, $m$ is TBA’s probability of sampling the data that is (relatively) the “most on-policy”; i.e., the data added at the most recent sync step.
Figure 2 visualizes our TBA implementation. It leverages overlapping training and search processes, and a buffer to hold trajectories. In particular, TBA uses a single Trainer node and one or more Searcher nodes that collect off-policy data into a shared replay buffer $\mathcal{D}_{\mathrm{global}}$. Here, a node is a computational resource sufficient to perform all the operations needed for training or search – it can be a collection of multiple GPUs or even a subset of 1 GPU. In our TBA experiments, a node is always 1 GPU: e.g., given 16 GPUs, we use 15 Searcher nodes and 1 Trainer node.
Separation of the Searcher and Trainer is highly desirable even in a 2 node (e.g. 2 GPU) cluster because LLM policy rollouts are costly sequential decoding procedures, and training requires only parallel likelihood evaluation of an entire sequence through a single forward pass. Thus, given an objective that can tolerate off-policy asynchronous online training, massive wall clock time speedups can be realized by continuously running training on 1 node without pausing for rollout generation. In the following subsections, we provide more details on our from-scratch TBA implementation.
@@ 4.1 Scaling data collection with Searcher nodes
TBA’s Searcher nodes each carry a local delayed copy $\pi_{\theta^{\prime}}$ of the Trainer policy $\pi_{\theta}$. To produce policy rollouts, queries $\mathbf{x}$ are sampled from a dataset, and the local policy generates a batch of $K$ responses $\mathbf{y}\sim\pi_{\theta^{\prime}}(\mathbf{y}|\mathbf{x})$ that are evaluated with the reward model $r_{\phi}(\mathbf{y};\mathbf{x})$. Like noukhovitch2024asynchronous, we use vLLM [kwon2023efficient] for faster generation. The $(\mathbf{x},\mathbf{y},r_{\phi}(\mathbf{y};\mathbf{x}))$ tuples are stashed in the Searcher’s local replay buffer $\mathcal{D}_{\text{local}}$. We also add to the stored tuple the step of the trainer when syncing last occurred, giving us a notion of how off-policy the generated data is – this can later be used by the Trainer to prioritize sampling from more recent generations.
Every $k$ optimization steps, search and training pause to pull each Searcher node’s local replay buffer $\mathcal{D}_{\text{local}}$ into the global replay buffer $\mathcal{D}_{\mathrm{global}}$, and to update the searcher’s local policy with the trainer’s. The global buffer maintains a list of all generated responses and rewards for each query.
A key motivation for asynchrony is its compatibility with rollout scaling for enhanced exploration or increased sequence lengths. For example, we generate $S>K$ samples for a given query – even when only updating the model using $K$ samples per query – to mitigate the lack of diversity caused by the fact that $K$ independent model rollouts are not guaranteed to produce $K$ unique sequences. Relatedly, future work could apply simple off-policy inference techniques in the Searcher nodes such as randomly sampling the softmax temperature, or using alternative decoding techniques like beam search. We expect such approaches to particularly aid solution discovery in sparse reward settings.
@@ 5.1 Tasks
  • Mathematical reasoning (MR): We study the GSM8K task which consists of grade-school level math problems and a binary reward based on exact match for the correct final answer [cobbe2021gsm8k]. We adopt the setup from kazemnejad2024vineppounlockingrlpotential, noukhovitch2024asynchronous, using an SFTed RhoMath-1B [lin2024rho] model as a base for RL post-training.
  • Preference fine-tuning (PFT): We consider the task of fine-tuning a language model with a reward function learned from human preference data. Specifically, we study the TL;DR summarization task where the goal is to write short summaries for reddit posts [stiennon2020learning]. Following noukhovitch2024asynchronous, we consider Pythia [biderman2023pythia] as the base model for the policy and the reward models, using the SFTed versions used by noukhovitch2024asynchronous.
  • Red-teaming (RT): We investigate automated red-teaming, another critical step for LLM post-training. The goal is to discover prompts that elicit harmful responses from a target model, as measured by a toxicity classifier. We follow the setup from lee2024learning, applying the same models: our smaller-scale experiments use GPT-2 [gpt2] as an attacker model, GPT-2 (instruction-tuned) as a victim model, and a RoBERTa-based toxicity classifier [vidgen2021lftw]; our larger-scale experiments use Llama-3.2-1B [metaAI2024llama3] as an attacker model, Llama-3.1-8B-Instruct as a victim model, and LlamaGuard-3-8B [metaAI2024llama3] for measuring toxicity, averaging over multiple responses from the victim model.
For all tasks, we study solely the RL post-training components of baselines’ workflows. In particular, we start with SFTed checkpoints, then perform RL. Notably, lee2024learning also include a maximum likelihood estimation training phase after RL that uses the buffer produced during RL, which boosts performance but is not investigated here. Using these baselines’ codes, we follow their setup and hyperparameters, with deviations noted in Appendix B. For MR and PFT, we implement TBA by augmenting the RLOO trainer of noukhovitch2024asynchronous with our distributed asynchronous RL framework and the TB objective (Equation 5). For RT, we implement TBA by augmenting the trajectory balance trainer of lee2024learning with our distributed asynchronous RL framework.
@@ 5.2 Metrics and baselines
MR metrics and baselines. We follow the evaluation setup of noukhovitch2024asynchronous, computing the pass@1 on the GSM8K [cobbe2021gsm8k] test dataset with greedy decoding. Prior work [kazemnejad2024vineppounlockingrlpotential, noukhovitch2024asynchronous] applies RL post-training with the GSM8K training set to the SFTed RhoMath-1B [lin2024rho] model, which initially obtains 40.3% accuracy on the test set. We compare TBA post-training to the methods used in these prior works: VinePPO [kazemnejad2024vineppounlockingrlpotential], Online-DPO [guo2024direct], (Async) PPO [schulman2017proximal], and (Proximal) RLOO [ahmadian2024back]. Additionally, we implement a GRPO [shao2024deepseekmath] baseline (see Appendix E.1 for details).
PFT metrics and baselines. We follow the evaluation setup of noukhovitch2024asynchronous, using the win-rate under a 6.7B “gold” reward model [huang2024n+] as the primary metric. We additionally report approximate KL distance—approximated by perplexity to match the evaluation of noukhovitch2024asynchronous—between the learned policy and the reference policy. Following noukhovitch2024asynchronous, we compare TBA with: Online-DPO [guo2024direct], (Async) PPO [schulman2017proximal], and (Proximal) RLOO [ahmadian2024back].
RT metrics and baselines. We follow lee2024learning, measuring the attack success rate on 1024 sampled prompts for a victim model. We also measure the diversity of these test-time generated prompts by computing the average pairwise cosine distance. We compare against: SFT, PPO [schulman2017proximal] (with the novelty reward from lee2024learning), REINFORCE [williams1992simple, sutton1999policy], RLOO [ahmadi-mahmudi-2023-revisiting], Online DPO [rafailov2024direct], GFlowNet (one-actor TB) [lee2024learning].
@@ 5.3 TBA redefines efficiency-performance Pareto frontiers
Figure 3: TBA scales search and improves RL efficiency on the PFT summarization task. All plotted points use 4xA100 GPUs, but TBA allocates 3 GPUs to search, and Online DPO allocates 1 GPU to search. TBA produces large-scale off-policy data that its trajectory balance objective can leverage, creating massive efficiency benefits. Online DPO baselines taken from noukhovitch2024asynchronous. Dashed and solid lines use 256 and 425 updates, respectively. Appendix B has details.
@@
When tested on established RL problems, an alternative possibility is that TBA will underperform due to its departures from conventional approaches: it is asynchronous, off-policy, reliant on the less-common trajectory balance objective, and makes updates using many responses per query. [^1] We compute the TBA loss with more samples per query than analogous approaches like RLOO (e.g., 20 samples versus 4) to reduce the variance of the gradient of the TB objective estimate (Equation 5) that we optimize. We use fewer queries per batch to keep the batch size small despite this response scaling. Indeed, noukhovitch2024asynchronous contemporaneously suggests potential limitations to asynchronous, off-policy RL for LLMs, finding that increasing off-policyness can harm performance metrics like win-rate or exacerbate policy deviations from the reference model. Further, hou2024does found limited benefits to scaling response generation from 4 to 8 (or 16) responses when using PPO.
@@
We study this question by computing Pareto frontiers for MR (Figure 1), for PFT (Figures 3 and 4 and Table 1), and for RT (Figure 5 and Table 6). See Appendix B for experimental details. Notably, TBA produces results on or beyond the Pareto frontiers of all three tasks at multiple model scales, consistent with our hypothesis that TBA can efficiently train LLMs on off-policy data.
Figure 4: TBA defines a new KL vs. win-rate Pareto frontier on the PFT summarization task. The baseline “Online DPO” frontier is created by increasing the degree of off-policyness, starting from on-policy Online DPO, results from [noukhovitch2024asynchronous]. The TBA frontier is created by altering the training steps, searcher count, and KL annealing schedule as described in Appendix B.
Table 1: TBA outperforms off-policy and on-policy baselines in the PFT task (Pythia 410M). Baselines from noukhovitch2024asynchronous. The final block uses 16 steps off-policy for baselines, and a sync period of 10 for TBA, which corresponds to use of data that is 15 steps off-policy on average.
Method
Steps/Sync Metric Online DPO (Async) PPO (Proximal) RLOO TBA
1 (on-policy) Perplexity/KL $\downarrow$ 1.13 1.14 1.13
Win Rate $\uparrow$ 0.85 0.86 0.82
2 (async) Perplexity/KL $\downarrow$ 1.13 1.14 1.13
Win Rate $\uparrow$ 0.85 0.84 0.83
$\approx 15$ (async) Perplexity/KL $\downarrow$ 1.12 1.10 1.10 1.13
Win Rate $\uparrow$ 0.82 0.76 0.77 0.86
@@
Speed is vastly improved with TBA training, which proceeds entirely asynchronously without training-bound or generation-bound processes – the only non-training time occurs briefly every $k$ steps. In the compute-matched MR experiments (Figure 1), TBA speeds up the training of the only method with comparable performance (VinePPO) by nearly $50\times$, while improving accuracy by 1.81.8% and speed by $1.5\times$ relative to the speed-optimized asynchronous DPO baseline [noukhovitch2024asynchronous]. In the compute-matched PFT experiments (Figure 3), TBA produces $\approx 5\times$ speedups over speed-optimized asynchronous DPO baselines [noukhovitch2024asynchronous], and it even outperforms baselines that train with on-policy data (Table 1). In the non-compute-matched automated red-teaming experiments (Table˜6), TBA gives $\approx 7\times$ speedups for GPT-2 and Llama 3.2 1B compared to the non-distributed, synchronous GFlowNet baseline [lee2024learning]. These results suggest that TBA is an effective, parallel, and scalable search framework for distributed learning, offering substantial speed-ups while remaining competitive with leading approaches.
@@ 5.4 Does off-policyness hurt performance?
The results in the prior section are perhaps surprising given evidence of off-policyness’s harmfulness to RL post-training [noukhovitch2024asynchronous]. Thus, we investigate the effect of off-policyness with TBA by studying its related hyperparameters, which we first review here. The fraction $m$ controls how often sampling attempts to approximate an “on-policy" distribution. When $m=1$, training occurs exclusively with samples added in the most recent sync step, while $m=0$ corresponds to selecting data without regard for how recently it was generated (e.g., using reward weighting). Additionally, recall that parameters and buffer data are shared between searchers and trainers every $k$ training steps: since TBA runs exploration and training in parallel, TBA trains off-policy even when $m=1$ and $k=1$. Specifically, with probability $m$, TBA selects a sample from the central replay buffer that is at most 2k12k-1 updates off-policy. With probability 1m1-m, TBA samples data produced by the model at any point, which can be as off-policy as the number of training steps.
@@
Importantly, the choice of reinforcement learning (RL) algorithm is crucial in determining how effectively we can leverage off-policy updates. TBA is a fully off-policy compatible algorithm, and as shown in Figure˜4, it significantly outperforms asynchronous Online DPO, even in the latter’s most on-policy setting. noukhovitch2024asynchronous identified Online DPO as the best-performing off-policy algorithm in their experiments (see Table 1), making it notable that TBA improves upon this.